# Dynamic Memory Compression: Retrofitting LLMs for Accelerated Inference

### v1

@@ Abstract
Transformers have emerged as the backbone of large language models (LLMs). However, generation remains inefficient due to the need to store in memory a cache of key–value representations for past tokens, whose size scales linearly with the input sequence length and batch size. As a solution, we propose Dynamic Memory Compression (DMC), a method for on-line key–value cache compression at inference time. Most importantly, the model learns to apply different compression rates in different heads and layers. We retrofit pre-trained LLMs such as Llama 2 (7B, 13B and 70B) into DMC Transformers, achieving up to ~3.7$\times$ throughput increase during auto-regressive inference on an NVIDIA H100 GPU. DMC is applied via continued pre-training on a negligible percentage of the original data without adding any extra parameters. We find that DMC preserves the original downstream performance with up to 4$\times$ cache compression, outperforming up-trained grouped-query attention (GQA). GQA and DMC can be even combined to obtain compounded gains. As a result DMC fits longer contexts and larger batches within any given memory budget.
@@
${}^{Q}$     ${}^{K}$niversity of Wrocław     ${}^{V}$niversity of Edinburgh
@@ 1 Introduction
Transformer Large Language Models (LLMs) are the state of the art in generative and conversational AI (Touvron et al., 2023; Jiang et al., 2023). Their deployment, however, is curtailed in part by their inefficiency. This is not only due to the quadratic complexity of attention layers (Bahdanau et al., 2014; Vaswani et al., 2017): during generation, Transformers store the keys and values of past tokens in memory to avoid recomputing them multiple times. Since this key–value (KV) cache grows linearly with the sequence length and batch size, generation with Transformers quickly becomes prohibitive due to the excessive memory load. This issue emerges even more clearly with long-context generation (e.g., in dialogues and stories) or when serving large numbers of user queries.
@@
A widespread solution to increase the memory efficiency of Transformers during inference is Grouped Query Attention (GQA; Ainslie et al., 2023; Shazeer, 2019), which uses a number of key and value heads inferior to the number of query heads through parameter sharing. As an alternative, the number of overall tokens in memory can be reduced through token merging (Zhang et al., 2018; Liu et al., 2018; Bolya et al., 2022) or token pruning (Anagnostidis et al., 2023; Kim & Cho, 2020). Nevertheless, these methods often pay the price of a severe degradation in downstream performance. On the other hand, hardware/IO-aware (Dao et al., 2022; Kwon et al., 2023) and sub-quadratic algorithms for attention (Beltagy et al., 2020; Choromanski et al., 2020) do not alleviate the memory load of the KV cache.
In our work, we aim to achieve a lossless compression of the KV cache of LLMs, thus retaining their performance while reducing their memory load. To this end, we propose Dynamic Memory Compression (DMC). As shown in Figure 1, during every time step, DMC decides whether to append the current key and value representations to the cache or to perform a weighted average of them with the top item on the cache. Note that memory grows sub-linearly in DMC, which therefore falls halfway between vanilla Transformers and state space language models (Fu et al., 2023; Gu & Dao, 2023), where memory is constant.
In our experiments, we equip pre-existing LLMs—such as Llama 2 (Touvron et al., 2023) 7B, 13B, and 70B—with DMC by retrofitting them on a negligible percentage of the original pre-training data (~2% for 2$\times$ compression and ~4% for 4$\times$ compression) and without adding any extra parameters to the original LLM. We evaluate our DMC models on a series of downstream tasks such as MMLU (Hendrycks et al., 2021) for factuality, QA datasets for common-sense reasoning, and HumanEval (Chen et al., 2021) for code. We find that DMC LLMs retain a downstream performance similar to the original LLM, whereas GQA incurs significant degradation. Finally, we show that DMC can be hybridized with GQA such that their compression rates are compounded. For Llama 2 70B, which is pre-trained with GQA 8×8$\times$8 ×, DMC 2×2$\times$2 × achieves a total compression of 16×16$\times$16 ×.
We verify that KV cache compression translates into more efficient generation in practice. We measure that DMC 4×4$\times$4 × increases the inference throughput between 340% and 370% for Llama 2 7B and 13B on NVIDIA H100 or A100 GPUs without loss in performance. In fact, it allows us to fit larger batches and longer sequences into a given memory budget. Finally, compression schemata learned by DMC provide insights into the internal structure of the LLMs, revealing a preference for compressing heads in higher layers.
@@ 2.1 Multi-Head Self-Attention
Let $X=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})\in\mathbb{R}^{n\times d}$ denote the input sequence of hidden states of a Transformer layer, where $n$ stands for the number of tokens in a sequence, and $d$ for the hidden state dimension. A Multi-head Self-Attention (MHSA) layer divides the embeddings into $n_{h}$ different heads. Afterwards, the self-attention process is applied to each head separately. This enables the model to focus on different parts of the input, capturing various types of relationships in the data. For each head $h$, different weight matrices $W_{q}^{h},W_{k}^{h},W_{v}^{h}\in\mathbb{R}^{d/n_{h}\times d/n_{h}}$ are used to project the input sequence into queries $Q^{h}$, keys $K^{h}$, and values $V^{h}$:
@@ 2.2 KV Caching During Inference
In a Transformer decoder, the generation of sequences is auto-regressive: each new token is predicted based on the previously generated ones. To avoid redundant computation, it is common to store the keys and values of previously computed tokens in the KV cache. For each time step $i$, only the keys and values for the current token $\mathbf{x}_{i}$ are computed whereas those for $\mathbf{x}_{<i}$ are retrieved from the cache. Thus for each head $h$:
@@ 3 Method: Dynamic Memory Compression
Inference with LLMs tends to be memory bound rather than compute bound, as we explain in Appendix A in detail. This means that latency scales linearly with the size of the KV cache. We tackle the problem of reducing the size of KV cache which brings two immediate benefits (Zhang et al., 2023): 1) it lowers the latency of auto-regressive generation, and 2) increases throughput by saving GPU memory and allowing for larger batch sizes or sequence lengths.
In this section we introduce Dynamic Memory Compression (DMC), a simple and inexpensive method for on-line compression of the KV cache at inference time. In what follows, we first describe the inference-time operation of DMC, which constitutes our end goal. Next, we illustrate how to teach a pre-trained LLM such behavior through short, continued pre-training.
@@ 3.1 Inference
Consider the forward pass of an attention layer during auto-regressive inference (Section 2.1). In a vanilla Transformer, at every time step $t$, both $\mathbf{k}_{t}$ and $\mathbf{v}_{t}$ are appended to the KV cache (Section 2.2). In DMC, on the other hand, the KV cache update procedure is different, as detailed in Algorithm 1. First, a decision variable $\alpha_{t}\in\{0,1\}$ and importance variable $\omega_{t}\in[0,1]$ are predicted. In order to avoid adding new parameters, we reuse the first neuron from $\mathbf{k}_{t}$ and $\mathbf{q}_{t}$, respectively, to extract the two scores. [^1] Note that this choice is arbitrary as we could use any two neurons from $\{\mathbf{q}_{t},\mathbf{k}_{t},\mathbf{v}_{t}\}$.
@@
[…]
Based on $\alpha_{t}$, a decision is made whether KV representations $\mathbf{k}_{t}$ and $\mathbf{v}_{t}$ are appended to the cache or accumulated with its last element (Figure 1). In particular, for accumulation we perform a weighted average based on a predicted importance scores $\omega$ for the current token and the running sum of importance scores $z_{t}$ for all the tokens since the last time step $\alpha=0$ was predicted.
In fact, the $\alpha$ variable effectively segments the input sequence: each decision determines if the current segment should continue ($\alpha=1$) or a new segment should be opened ($\alpha=0$). As a result, after the update, the cache length for DMC is $l=\sum_{i=1}^{t}(1-\alpha_{i})\leq t$, whereas in vanilla Transformers it is always $l=t$. In what follows, we will refer to the ratio $t/l$ between the length $t$ of an uncompressed cache and the compressed length $l$ as the Compression Ratio (CR).
Finally, multi-head self-attention is calculated similarly to vanilla Transformer using KV cache sequences, with the exception that KV sequences for different heads might have different lengths. Algorithm 1 is applied to every MHSA layer and head independently.
@@ 3.2 Training
The DMC inference algorithm switches between accumulating and appending tokens to the KV cache. In order to endow LLMs with DMC, we continue pre-training them on a negligible amount of their pre-training data, gradually increasing the compression rate towards a target. However, this poses serious challenges. First, we opt for end-to-end learning via gradient descent and continuous relaxation of the decision variables. As a result, we have to define an operation for KV cache updating when $10<\alpha<$ 1, resulting in partly aggregated, partly accumulated key and value states. Second, to avoid training–inference mismatch, we must simulate the DMC behavior at inference time while parallelizing training across a sequence of tokens: as a consequence the length of $K$ and $V$ is not reduced through compression during training; rather, all intermediate states of keys and values are explicitly kept in memory and an auxiliary (gradually discretized) mask regulates query interactions. We elaborate on our solutions to these challenges below.
@@
The decision whether to accumulate or append at inference time is a discrete one; however, rounding $\mathrm{sigmoid}(\mathbf{k}[0])$ to the nearest integer for training would result in a non-differentiable operation with zero gradients. Hence, we resort to stochastic reparametrization of the decision variable during training
$\alpha_{t}\sim\text{Gumbel-sigmoid}(\mathbf{k}[0]-c,\tau)\in[0,1],$ (8)
where $\tau$ is the temperature [^2] Low temperatures sharpen $\alpha_{t}$ into almost-discrete values which accurately mimics inference behavior. and $c$ is a constant subtracted so that every $\alpha=0$ at training step 00. This ensures that DMC initially performs no compression and starts the training behaving just like the vanilla Transformer.
@@
As we relax our discrete decisions, we now must define a mechanism to update the KV cache that generalizes Algorithm 1 to continuous $\alpha$. Hence, we define partially accumulated states for $\alpha\in[0,1]$ as:
@@
Aside from key and value computations shown in Equation 9, the rest of the forward pass can be performed in parallel for all tokens in the sequence. Nonetheless, this creates a mismatch between training and evaluation, since all intermediate states of keys and values are accessible during self-attention.
To illustrate this issue, consider the example of a KV cache during DMC inference shown in Figure 2 for the sequence of decision scores $\alpha_{1:5}=(1,1,0,1,0)$ (importance scores $\omega$ have been omitted for clarity). The last element of the KV cache changes at every time step. In order to properly simulate the inference-time evolution of the KV cache during training, we keep all unrolled intermediate KV cache items. In lieu of an auto-regressive ‘causal’ mask, however, we use an additive mask based on the sequence of $\alpha$ values to modify the attention scores $a^{h}_{ij}$ from Equation 4, which is shown in Figure 3. During training, the $\alpha$ values 1) naturally converge towards 0 or 1 as the model strives to satisfy the language modeling criterion and reduce uncertainty; 2) are intentionally pushed to almost-discrete states by the Gumbel noise and low temperature setting. Such binarization of $\alpha$ values significantly impacts the attention scores - it strengthens the queries interactions with last elements of every key–value segment, and weakens those with the intermediate elements, which are discarded during inference.333 See Appendix F for details on the mask implementation. In fact, when $\alpha\in\{0,1\}$, the matrix is filled with either 00 or $\infty$- ∞ values, and exactly corresponds to the inference time query-to-key attendance pattern.
@@
Figure 3: Additive mask applied during training to the the normalized attention scores to block queries from attending intermediate KV states (as well as future states).
@@
The model is incentivized to compress the KV cache to a certain CR, and thus increase the predicted $\alpha$ values. Instead of matching the desired rate for each append-or-accumulate decision $\alpha$, we calculate a global one-sided loss as the difference between the expected sum of KV tokens across all layers $l$, heads $h$ and time steps $t$ under the desired Compression Rate CR and the sum of all decisions, normalized by $(n_{l}n_{h}n)$:
$\ell_{\text{CR}}=\frac{1}{n_{l}n_{h}n}*max(0,\frac{n_{l}n_{h}n}{\text{CR}}-\sum_{l=1}^{n_{l}}\sum_{h=1}^{n_{h}}\sum_{t=1}^{n}\alpha_{lht}).$ (10)
It is added to the language modeling loss term $\ell_{\text{LM}}=-\sum_{t=1}^{n}\log p_{\boldsymbol{\theta}}(\mathbf{x}_{t}\mid\mathbf{x}_{<t})$, with the final objective of the training being:
@@
Importantly, the training procedure is designed for slowly ramping up the desired compression rate and stopping at will. This enables producing a series of models with different compression rates within a single run. It follows that all hyperparameters, like Gumbel-sigmoid sampling temperature and learning rate, are not decayed and remain constant throughout training.

### v2

@@ Abstract
Transformers have emerged as the backbone of large language models (LLMs). However, generation remains inefficient due to the need to store in memory a cache of key–value representations for past tokens, whose size scales linearly with the input sequence length and batch size. As a solution, we propose Dynamic Memory Compression (DMC), a method for online key–value cache compression at inference time. Most importantly, the model learns to apply different compression ratios in different heads and layers. We retrofit pre-trained LLMs such as Llama 2 (7B, 13B, and 70B) into DMC Transformers, achieving up to 7$\times$ throughput increase during auto-regressive inference on an NVIDIA H100 GPU. DMC is applied via continued pre-training on a negligible percentage of the original data without adding any extra parameters. DMC preserves the original downstream performance with up to 4$\times$ cache compression, outperforming up-trained grouped-query attention (GQA) and key–value eviction policies (H2O, TOVA). GQA and DMC can be even combined to obtain compounded gains. Hence, DMC can serve as a drop-in replacement for KV caching in existing LLMs to fit longer contexts and larger batches within any given memory budget.
@@
QNVIDIA     KUniversity of Wrocław     VUniversity of Edinburgh
@@ 1 Introduction
Transformer Large Language Models (LLMs) are the state of the art in generative and conversational AI (Touvron et al., 2023; Jiang et al., 2023). Their deployment, however, is curtailed in part by their inefficiency. This is not only due to the quadratic complexity of attention layers (Bahdanau et al., 2015; Vaswani et al., 2017): during generation, Transformers store the keys and values of past tokens in memory to avoid recomputing them multiple times. Since this key–value (KV) cache grows linearly with the sequence length and batch size, generation with Transformers quickly becomes prohibitive due to the excessive memory load. This issue emerges even more clearly with long-context generation (e.g., in dialogues and stories) or when serving large numbers of user queries.
@@
A widespread solution to increase the memory efficiency of Transformers during inference is Grouped Query Attention (GQA; Ainslie et al., 2023; Shazeer, 2019), where the number of key and value heads is reduced by sharing each of them among multiple query heads. Alternatively, the length of the KV cache can be shortened through token merging (Zhang et al., 2018; Liu et al., 2018; Bolya et al., 2023) or cache eviction policies (Zhang et al., 2023; Oren et al., 2024). Nevertheless, these methods often pay the price of a severe degradation in downstream performance. On the other hand, hardware/IO-aware (Dao et al., 2022; Kwon et al., 2023) and sub-quadratic algorithms for attention (Beltagy et al., 2020; Choromanski et al., 2021) do not alleviate the memory load of the KV cache.
In our work, we aim to adaptively compress the KV cache of LLMs, retaining their performance while reducing their memory load. To this end, we propose Dynamic Memory Compression (DMC). As shown in Figure 1, during every time step, DMC decides whether to append the current key and value representations to the cache or to perform a weighted average of them with the top-most item in the cache. Note that, for an appropriate prior over compression decisions, the memory can grow sub-linearly in DMC, which therefore falls halfway between vanilla Transformers and state space language models (Fu et al., 2023; Gu & Dao, 2023), where memory size is constant.
In our experiments, we equip pre-existing LLMs—such as Llama 2 (Touvron et al., 2023) 7B, 13B, and 70B—with DMC by retrofitting them on a negligible percentage of the original pre-training data (~2% for 2$\times$ compression, and ~8% for 8$\times$ compression) and without adding any extra parameters to the original LLM. We evaluate our DMC models on a series of downstream tasks such as MMLU (Hendrycks et al., 2021) for factuality, QA datasets for common-sense reasoning, and HumanEval (Chen et al., 2021) for code. We find that DMC LLMs retain a downstream performance similar to the original LLM, whereas baselines—such as GQA, H2O, and TOVA—incur significant degradation at high compression ratios. Finally, we show that DMC can be hybridized with GQA such that their compression ratios are compounded. For Llama 2 70B, which is pre-trained with GQA 8×8$\times$8 ×, DMC 2×2$\times$2 × achieves a total compression of 16×16$\times$16 ×.
We verify that KV cache compression translates into more efficient generation in practice. We measure that DMC 4×4$\times$4 × increases the inference throughput between 350% and 390% for Llama 2 7B and 13B on NVIDIA H100 or A100 GPUs without loss in performance. For higher compression, such as DMC 8×8$\times$8 ×, we observe inference throughput gains up to 700% with ~5% relative MMLU performance drop for Llama 2 7B and 13B. In fact, DMC allows us to fit larger batches and longer sequences into a given memory budget. Finally, compression schemata learned by DMC provide insights into the internal structure of the LLMs, revealing a preference for compressing heads in higher layers.
@@ 2.1 Multi-Head Self-Attention
Let $X=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})\in\mathbb{R}^{n\times d}$ denote the input sequence of hidden states of a Transformer layer, where $n$ stands for the number of tokens in a sequence, and $d$ for the hidden state dimension. A Multi-Head Self-Attention (MHSA) layer divides the embeddings into $n_{h}$ different heads. Afterward, the self-attention process is applied to each head separately. This enables the model to focus on different parts of the input, capturing various types of relationships in the data. For each head $h$, different weight matrices $W_{q}^{h},W_{k}^{h},W_{v}^{h}\in\mathbb{R}^{d/n_{h}\times d/n_{h}}$ are used to project the input sequence into queries $Q^{h}$, keys $K^{h}$, and values $V^{h}$:
@@ 2.2 KV Caching During Inference
In a Transformer decoder, the generation of sequences is auto-regressive: each new token is predicted based on the previously generated ones. To avoid redundant computation, it is common practice to store the keys and values of previously computed tokens in the KV cache. For each time step $i$, only the keys and values for the current token $\mathbf{x}_{i}$ are computed whereas those for $\mathbf{x}_{<i}$ are retrieved from the cache. Thus for each head $h$:
@@
2.3 Memory-Bound and Compute-Bound Operations
Every operation performed with a GPU accelerator, such as General Matrix Multiply (GEMM), is either memory-bound or compute-bound. In the former case, the overall runtime is dominated by high bandwidth memory (HBM) access, while in the latter by the actual computations. Auto-regressive generation with Transformer LLMs, where the sequence length for every forward pass is $n=1$, tends to be memory-bound rather than compute-bound. The vast majority of a forward pass is spent either processing linear layers (in MHSA, Feed-Forward, and output vocabulary projection) or calculating attention scores and outputs from Equation 4. For linear layers, the ratio of FLOPS to memory accesses improves as the batch size increases, and more FLOPS are performed with the set of layer weights retrieved from the HBM. Eventually, with a large enough batch size, linear layers become compute-bound. On the other hand, for the calculation of Equation 4 inside MHSA layers during auto-regressive inference, the ratio of FLOPS to input size remains constant, and MHSA layers are memory-bound regardless of the batch size. It follows that for those layers, latency scales linearly with the size of the KV cache.
@@ 3 Method: Dynamic Memory Compression
In this work, we tackle the problem of reducing the size of the KV cache, which brings two immediate benefits (Zhang et al., 2023): 1) it lowers the latency of auto-regressive generation, and 2) improves GPU utilization by allowing for larger batch sizes and sequence lengths, leading to increased throughput. In this section, we introduce Dynamic Memory Compression (DMC), a simple and inexpensive method for online compression of the KV cache at inference time. In what follows, we first describe the inference-time operation of DMC, which constitutes our end goal. Next, we illustrate how to teach a pre-trained LLM such behavior through short, continued pre-training.
@@ 3.1 Inference
Consider the forward pass of an attention layer during auto-regressive inference (Section 2.1). In a vanilla Transformer, at every time step $t$, both $\mathbf{k}_{t}$ and $\mathbf{v}_{t}$ are appended to the KV cache (Section 2.2). In DMC, on the other hand, the KV cache update procedure is different, as detailed in Algorithm 1. First, a decision variable $\alpha_{t}\in\{0,1\}$ and importance variable $\omega_{t}\in[0,1]$ are predicted. In order to avoid adding new parameters, we reuse the first neuron from $\mathbf{k}_{t}$ and $\mathbf{q}_{t}$, respectively, to extract the two scores: [^1] Note that this choice is arbitrary as we could use any two neurons from $\{\mathbf{q}_{t},\mathbf{k}_{t},\mathbf{v}_{t}\}$.
$\alpha_{t}\mkern 1.5mu{=}\mkern 1.5mu\lfloor\text{sigmoid}(\mathbf{k}_{t}[0])\rceil,\ \omega_{t}\mkern 1.5mu{=}\mkern 1.5mu\text{sigmoid}(\mathbf{q}_{t}[0]).$ (8)
where sigmoid normalizes both scores into a range $[0,1]$. At inference time, the decision score $\alpha_{t}$ is obtained by rounding to the closest integer to make this variable binary.
@@
0:  $K,V,\mathbf{q}_{t},\mathbf{k}_{t},\mathbf{v}_{t}$
1:  $\alpha_{t}\leftarrow\text{round}(\text{sigmoid}(\mathbf{k}_{t}[0]))$
2:  $\omega_{t}\leftarrow\text{sigmoid}(\mathbf{q}_{t}[0])$
3:  $\mathbf{k}_{t}[0],\mathbf{q}_{t}[0]\leftarrow 0,0$
4:  if $\alpha_{t}=1$ then  $\triangleright$ accumulate
5:     $z_{t}\leftarrow z_{t-1}+\omega_{t}$
6:     $K\leftarrow[K_{1:l-1},(K_{l}z_{t-1}+\mathbf{k}_{t}\omega_{t})/z_{t}]$
7:     $V\leftarrow[V_{1:l-1},(V_{l}z_{t-1}+\mathbf{v}_{t}\omega_{t})/z_{t}]$
8:  else  $\triangleright$ append
9:     $z_{t}\leftarrow\omega_{t}$
10:     $K\leftarrow[K_{1:l},\mathbf{k}_{t}]$
11:     $V\leftarrow[V_{1:l},\mathbf{v}_{t}]$
12:  end if
13:  return $K,V,\mathbf{q}_{t},\mathbf{k}_{t}$
Based on $\alpha_{t}$, a decision is made whether KV representations $\mathbf{k}_{t}$ and $\mathbf{v}_{t}$ are appended to the cache or accumulated with its last element (Figure 1). In particular, for accumulation, we perform a weighted average based on the predicted importance score $\omega_{t}$ for the current token and the running sum of importance scores $z_{t}$ for all the tokens since the last time step $\alpha=0$ was predicted.
In effect, the $\alpha$ variable segments the input sequence: each decision determines if the current segment should continue ($\alpha=1$) or a new segment should be opened ($\alpha=0$). As a result, after the update, the cache length for DMC is $l=\sum_{t=1}^{n}(1-\alpha_{t})\leq n$, whereas in vanilla Transformers it is always $l=n$. In what follows, we will refer to the ratio $n/l$ between the length $n$ of an uncompressed cache and the compressed length $l$ as the Compression Ratio (CR).
Finally, multi-head self-attention is calculated similarly to vanilla Transformers but using the compressed KV cache sequences. Algorithm 1 is applied to every MHSA layer and head independently. Hence, the corresponding KV sequences might have different lengths. Note that Algorithm 1 can be implemented efficiently without the if-then-else statement conditioned on $\alpha_{t}$, instead multiplying the previous ${\mathbf{k}}_{i}$, ${\mathbf{v}}_{i}$ and $z_{i}$ by $\alpha_{t}$ like in Equation 10.
@@ 3.2 Training
The DMC inference algorithm switches between accumulating and appending tokens to the KV cache. In order to endow LLMs with DMC, we continue pre-training them on a negligible amount of their pre-training data, gradually increasing the compression ratio towards a target. However, this poses serious challenges. First, we opt for end-to-end learning via gradient descent, which requires a continuous relaxation of the decision variables. As a result, we have to define an operation for KV cache updating when $10<\alpha<$ 1, resulting in partly aggregated, partly accumulated key and value states. Second, to avoid training–inference mismatch, we must simulate the DMC behavior at inference time while parallelizing training across a sequence of tokens. As a consequence the length of $K$ and $V$ is not reduced through compression during training; rather, all intermediate states of keys and values are explicitly kept in the sequence and an auxiliary, gradually discretized mask regulates interactions between queries and keys. We elaborate on our solutions to these challenges below.
@@
The decision whether to accumulate or append at inference time is a discrete one; however, rounding $\alpha_{t}=\mathrm{sigmoid}(\mathbf{k}_{t}[0])$ to the nearest integer for training would result in an operation with zero or undefined gradients. Hence, we resort to stochastic reparametrization of the decision variable during training
$\alpha_{t}\sim\text{Gumbel-sigmoid}(\mathbf{k}_{t}[0]-c,\tau)\in[0,1],$ (9)
where $\tau$ is the temperature [^2] Low temperatures sharpen $\alpha_{t}$ into almost-discrete values which accurately mimics inference behavior. and $c$ is a constant subtracted so that every $\alpha\approx 0$ at training step 00. This ensures that DMC initially performs no compression and starts the training behaving just like the vanilla Transformer.
@@
As we relax our discrete decisions, we must now define a mechanism to update the KV cache that generalizes Algorithm 1 to continuous $\alpha$. Hence, we define partially accumulated states for $\alpha\in[0,1]$ as:
@@
Aside from key and value computations shown in Equation 10, the rest of the forward pass can be performed in parallel for all tokens in the sequence. Nonetheless, this creates a mismatch between training and evaluation, since during training all intermediate states of keys and values are accessible in self-attention.
To illustrate this issue, consider the example of a KV cache during DMC inference shown in Figure 2 for the sequence of decision scores $\alpha_{1:5}=(1,1,0,1,0)$ (importance scores $\omega$ have been omitted for clarity). Because the last element of the KV cache changes in-place at every time step, the future elements should not have access to its old contents (Figure 2, right). In order to properly simulate this inference-time evolution of the KV cache during training, we keep all unrolled intermediate KV cache items. However, in lieu of an auto-regressive ‘causal’ mask, we use an additive mask based on the sequence of $\alpha$ values to modify the attention scores $a^{h}_{ij}$ from Equation 4, which is shown in Figure 3. As $\alpha$ values are increasingly pushed to almost-discrete states by the Gumbel estimator and the low-temperature setting, this strengthens the interactions of queries with the last state of every key–value segment (i.e., the only one accessible during inference), and weakens those with the previous states, which are discarded during inference.333 See Appendix G for details on the mask implementation. In fact, when $\alpha\in\{0,1\}$, the matrix is filled with either 00 or $\infty$- ∞ values, and exactly corresponds to the inference-time query-to-key attendance pattern.
@@
Figure 3: Additive mask applied during training to the normalized attention scores in order to block queries from attending intermediate KV states (as well as future states).
@@
The model is incentivized to compress the KV cache to a certain CR, and thus increase the predicted $\alpha$ values. Instead of matching the desired rate for each append-or-accumulate decision $\alpha$, we calculate a global one-sided loss as the difference between the sum of all decisions and the expected sum of KV tokens across all layers $l$, heads $h$ and time steps $t$ under the desired Compression Ratio (CR), normalized by $(n_{l}n_{h}n)$:
$\ell_{\text{CR}}=\frac{1}{n_{l}n_{h}n}\text{max}\left(0,\sum_{l=1}^{n_{l}}\sum_{h=1}^{n_{h}}\sum_{t=1}^{n}(1-\alpha_{lht})-\frac{n_{l}n_{h}n}{\text{CR}}\right).$ (11)
The $\ell_{\text{CR}}$ loss is added to the language modeling loss term $\ell_{\text{LM}}=-\sum_{t=1}^{n}\log p_{\boldsymbol{\theta}}(x_{t}\mid\mathbf{x}_{<t})$, with the final objective of the training being:
@@
Importantly, the training procedure is designed for slowly ramping up the target CR and saving ready-to-use DMC checkpoints along the way. This is possible because all hyperparameters, like Gumbel-sigmoid sampling temperature and learning rate, are not decayed and remain constant throughout training. A practical use case of this DMC property is to produce a series of DMC checkpoints with different CRs within a single run, and then choose the one with the desired efficiency–performance trade-off.